OpenStudy (anonymous):

What is the exact volume of a sphere that has a radius of 8.4 in.? [ ? ] (pi) in3

OpenStudy (anonymous):

@sleepyjess @iGreen

OpenStudy (sleepyjess):

First what is the formula for volume of a sphere?

OpenStudy (anonymous):

\[V=\frac{ 4 }{ 3 }\pi r^{2}\]

OpenStudy (sleepyjess):

Close, it should be \(\dfrac 43 \pi r^3\)

OpenStudy (anonymous):

sorry, thats what i meant

OpenStudy (sleepyjess):

Ok :)

OpenStudy (sleepyjess):

so just substitute in r

OpenStudy (anonymous):

754.654

OpenStudy (sleepyjess):

How did you get that?

OpenStudy (sleepyjess):

\(\dfrac 43 * 3.14 * 8.4^3\)

OpenStudy (anonymous):

2481.4540799999958

OpenStudy (sleepyjess):

I'm getting 2481.4508, but my calculator doesn't go as many decimals out as your it looks like

OpenStudy (anonymous):

ya

OpenStudy (sleepyjess):

They don't tell you to round to a certain decimal place?

OpenStudy (sleepyjess):

OOOHHHH exact volume, leave it in terms of \(\pi\)

OpenStudy (anonymous):

ya not 3.14

OpenStudy (sleepyjess):

yep, sorry

OpenStudy (anonymous):

your fine

OpenStudy (anonymous):

what do we do now?

OpenStudy (anonymous):

@sleepyjess ???

OpenStudy (anonymous):

@iGreen @BloomLocke367

OpenStudy (bloomlocke367):

what do you need help with?

OpenStudy (anonymous):

the question

OpenStudy (anonymous):

he left with out finishing!

OpenStudy (bloomlocke367):

don't multiply out the pi... you'll have the pi symbol in the answer

OpenStudy (anonymous):

OR SHE

OpenStudy (anonymous):

ok

OpenStudy (bloomlocke367):

so.. what's 8.4^3?

OpenStudy (bloomlocke367):

iGreen, don't you dare steal this one XD

OpenStudy (igreen):

\(V = \dfrac{4}{3} r^3\) Plug in 8.4 for 'r', this will give you the answer in terms of pi.

OpenStudy (igreen):

xD

OpenStudy (igreen):

I was tagged :P

OpenStudy (bloomlocke367):

I got here first XD

OpenStudy (anonymous):

790.272

OpenStudy (anonymous):

the more the merrier!

OpenStudy (igreen):

Correct! :P

OpenStudy (bloomlocke367):

Yep, good job!

OpenStudy (igreen):

\(790.272 \pi~in^3\)

OpenStudy (anonymous):

ok thx!

OpenStudy (sleepyjess):

Sorry, I had to go do something :(